# Online Smoothed Demand Management

### v1

@@ 1 Introduction
The proliferation of renewable energy sources and grid-scale energy storage, the growth of power-hungry AI data centers, and the implementation of dynamic pricing schemes like time-of-use tariffs are fundamentally reshaping the landscape of how energy and compute infrastructures interact. These changes are prompting a paradigm shift in how systems interact with power grids, shifting from a model of passive consumption to one of active, intelligent demand management. The core challenge in demand management from an energy perspective is to better align the consumption of energy (e.g., from the grid) with external signals such as time-of-use prices, while meeting certain performance (e.g,. service-level agreement (SLA)) objectives on the demand side. Concurrently, there is a need to consider the “smoothness” of these decisions: recent reports suggest that data centers’ sudden large fluctuations in energy use are impacting power quality in U.S. grids Bloomberg:24, and sudden changes in load are known to cause cascading failures and even blackouts on the grid, such as the 2025 Iberian Peninsula blackout Pons:2025. This prompts an additional objective of smooth decisions (e.g., gradual ramping up and down).
@@
Many techniques have been developed for this class of problems, with solution designs that are usually tailored to specific applications such as energy storage management Mamun:16:multi-objective-DR-battery and HVAC control Kou:21:HVAC, among others. These techniques often rely on machine learning somewhere in the pipeline, e.g., to provide predictions for a model predictive controller Serale:18:MPC-HVAC; Morstyn:17:MPC-battery, inform a day-ahead provisioning problem Yang:25:CarbonAwareHVAC, or to directly make decisions (e.g., using a reinforcement learning agent) Azuatalam:20. While these problem-specific solutions show promise, they are seldom deployed in practice. One reason is that environments such as electricity markets are inherently nonstationary, making it difficult to develop heuristics that deal well with distribution shift Levin:22:Texas21 or come with rigorous performance guarantees. Another reason is that these approaches are typically developed with a single end application in mind, making generalization to other settings difficult.
@@
On the other side of the spectrum, there is a rich body of work in online algorithms Yang:20; Lechowicz:24; ElYaniv:01; Yang:24; Chen:22:InBev on related problems that are sufficiently abstract to model many different applications, and have been studied under frameworks such as competitive analysis, which provides rigorous performance guarantees. Despite this, however, there remains a gap between the existing models studied in the online algorithms literature and the complexity of real-world demand management problems – as of yet, no models exist that fully capture the moving pieces necessary to study the type of problem that arises in, e.g., a grid-integrated data center with local energy storage.
@@
Competitive (i.e., worst-case) guarantees are often overly pessimistic in practice, since real-world problem instances are not adversarial. In recent years, the area of learning-augmented algorithms Lykouris:18; Purohit:18; Mitzenmacher:22:ALPS has expanded the online algorithms literature with techniques that can take additional information as input (e.g., predictions about the instance) to improve average-case performance while retaining worst-case guarantees for robustness against nonstationarities. While this paradigm has successfully been applied to problems that share similarities with OSDM Lechowicz:24; SunLee:21, traditional learning-augmented designs make strong assumptions about the availability and learnability of suitable predictions, which face challenges in OSDM due to the multiple kinds and nonstationary sources of uncertainty (e.g., demand arrivals, deadlines, and market prices). Thus, we also ask:
Can we design data-driven algorithms for OSDM that directly learn from data (e.g., historical instances) to significantly improve performance while preserving worst-case guarantees?
1.1 Contributions
@@
In § 4, we provide the first competitive algorithm for OSDM, which we call PAAD (Partitioned Accounting & Aggregated Decisions, see Algorithm 2). PAAD introduces a driver abstraction that splits OSDM into subproblems that can be solved independently and aggregated to form a globally competitive solution (Theorem 4.1). PAAD reflects, to the best of our knowledge, the most complex extension of the online pseudo-cost minimization algorithm design paradigm SunZeynali:20; we employ multiple kinds and instances of threshold functions to decompose and accommodate the significant heterogeneity of the OSDM setting. PAAD achieves a worst-case competitive ratio that is sublinear in $\nicefrac{{p_{\max}}}{{p_{\min}}}$ (the ratio of the procurement price bounds) and linear in other parameters of the problem setting such as the switching and delivery cost magnitudes – in § 5, we give lower bounds that imply PAAD achieves the best-possible competitive ratio for OSDM under our assumptions.
@@
To improve average-case performance and reduce the pessimism of our approach, in § 6 we present a learning framework for OSDM called PALD (Partitioned Accounting & Learned Decisions). Compared to existing work on learning-augmented algorithms, the OSDM setting brings unique challenges that make previously-considered techniques for integrating predictions impractical or ineffective. To address this challenge, PALD proposes a novel end-to-end differentiable learning framework that directly predicts the best algorithm parameters (e.g., threshold functions) that minimize PALD’s cost on historical instances, while ensuring that the learned parameters satisfy a robustness certificate (see Theorem 6.1) that guarantees a bounded worst-case competitive ratio under adversarial instances, serving as a safeguard against nonstationarity. This framework will likely be of independent interest as a more flexible and expressive methodology to learn high-performing and robust algorithms across problem domains.
@@
Finally, in § 7, we present a case study of OSDM in a simulation of a grid-integrated data center with co-located battery storage, using real electricity price data and production Alibaba traces Alibaba:18. We show that PAAD solves the OSDM problem effectively in practice, and that PALD’s end-to-end learning capability significantly improves performance compared to PAAD using data-driven insights.
1.2 Related Work
@@
Our study of OSDM is informed by and builds upon related work in online algorithms, learning-augmented algorithms, and relevant application areas. We briefly summarize these below.
@@
Applications of Demand Management and Related Ideas. OSDM-like problems arise across multiple domains – below, we review some heuristic and control-theoretic approaches that have been proposed to solve these. Given forecasts (e.g., prices), a common approach is to solve an offline “day-ahead” optimization problem, as applied in battery arbitrage MohsenianRad:16:day-ahead-battery; Zhang:24:day-ahead-battery; Faraji:20:day-ahead-prosumer, microgrids Silva:20:day-ahead-microgrid; Tandukar:18:day-ahead-VPP, HVAC Yang:25:CarbonAwareHVAC; Zeng:23:day-ahead-AC, and sustainable computing radovanovic2022carbon; Li:17:battery-DC. In online settings, a frequently-applied technique is model predictive control (MPC), which uses forecasts to solve a finite-horizon optimization problem at each time step. MPC has been applied to battery management Blonsky:20:MPC-bat-management; chehade:2024:MPC-bat-management; Sage:24:RL-battery-storage; Morstyn:17:MPC-battery, building systems Afram:15:MPC-HVAC; Serale:18:MPC-HVAC; Blum:22:MPC-HVAC; Mariano:21:BMS, microgrids MorenoCastro:23:microgrids; Lopez:21:MPC-hybridpp, and data centers Wang:23:hierarchical; Lazic:18:MPC-BMS; Mirhoseininejad:21:DC. Select works consider approaches that do not require predictions such as Lyapunov-based control Urgaonkar:11:lyapunov-power-cost-management; Guo:13:Lyapunov-battery-DC; Yu:15:Lyapunov-battery-DC; Stai:21:BatteryStorage; Sun:23:Lyapunov-battery-DC, ADMM for distributed storage RigoMariani:22:ADMM, feedback-based demand response Mamun:16:multi-objective-DR-battery, and stochastic optimization Nasiriani:17:stochastic-peak-shaving; Bedi:18. These approaches are generally developed with a single application in mind, limiting their generalizability, and do not typically come with theoretical guarantees on performance, particularly in cases where predictions or stochastic models are used.
@@
Related Online Algorithms Problems. In the literature there are a few online problems that share structure with OSDM. Of these, perhaps the closest related problem is online linear inventory management (OLIMYang:20, which models the problem faced by an operator (e.g., a data center) with an inventory (e.g., battery) that must meet a demand at each time step, while minimizing their procurement cost subject to time-varying prices. Compared to OLIM, OSDM additionally models a delivery cost to serve demand (i.e., separate from procurement), smoothing penalties, and a mixture of base (inelastic) and flexible demand. A recent model called online conversion with switching costs (OCSLechowicz:24 has modeled the problem faced by an operator procuring energy (e.g., an EV charger) that must meet a single flexible demand before a deadline (e.g., charging an EV by the end of the day), paying a switching cost when decisions change. OSDM generalizes this by including inflexible (i.e., base) demand constraints, delivery costs, and multiple flexible units with heterogeneous deadlines. Other formulations address energy storage without time-varying procurement prices Kim:16; Mo:21 or microgrid scheduling with variable prices but no storage Lu:13; Hajiesmaili:16; Zhang:2018:microgrid.
@@
Each component dynamic of OSDM has been studied in isolation: procurement and flexible demand via online knapsack and search Zhou:08; ElYaniv:01; Fung:21, per-step demand via online matching Mehta:07; Huang:24, and smoothness penalities via metrical task systems Borodin:92; FriedmanLinial:93 However, the combination of these in OSDM introduces new technical challenges in the design of robust and efficient algorithms that preclude the direct application of existing results (see, e.g., § 3). Online resource allocation problems jaillet2011online; Devanur:19; Wang:17:ORA; Jiao:17:ORA; Klimm:19; Zhang:17:ORA; Balseiro:23 are also related, though they usually assume a fixed (non-replenishable) resource. Some variants consider inventories Larsen:10; Cheung:21; Chen:22:InBev; He:22; HIHAT:23; Yang:24, but typically treat either demand or inventory evolution as exogenous, whereas in OSDM both are decisions.
2 Problem Formulation
@@
2.1 Online Smoothed Demand Management (OSDM)
Problem Statement. Consider an operator who makes decisions over a discrete time horizon $t\in\{1,\dots,T\}$. At each time step $t$, several inputs are revealed online. The first of these is a market price for energy, denoted by $p_{t}\geq 0$. Furthermore, two types of demand arrive: a base (inflexible) demand $b_{t}\geq 0$ that must be satisfied immediately, and a flexible demand $f_{t}\geq 0$ that can be deferred until a specified deadline $\Delta_{t}>t$. Further, the operator can control the state of a local energy storage unit (e.g., a battery) with a maximum capacity of $S\geq 0$. We denote the state of charge at time step $t$ as $s_{t}$, where the initial storage state is empty (i.e., $s_{0}=0$). [^1] The assumption that the initial storage state is $s_{0}=0$ is without loss of generality – our results hold when $s_{0}>0$.
@@ 2.1 Online Smoothed Demand Management (OSDM)
$[\texttt{OSDM}]\min_{\{x_{t},z_{t}\}_{t\in[T]}}$ $\sum\nolimits_{t=1}^{T}\Big[\underbrace{p_{t}x_{t}}_{\text{purchase cost}}+\underbrace{\mathcal{D}(z_{t},s_{t-1},p_{t})}_{\text{delivery cost}}\Big]+\sum\nolimits_{t=1}^{T+1}\Big[\underbrace{\mathcal{S}(x_{t},x_{t-1})+\delta|z_{t}-z_{t-1}|}_{\text{smoothing penalty}}\Big],$ (1)
s.t. $s_{t}=s_{t-1}+x_{t}-z_{t},\text{ with }0\leq s_{t}\leq S,$
$z_{t}-s_{t-1}\leq x_{t}\leq z_{t}+(S-s_{t-1}),$
@@
$\sum\nolimits_{\tau=1}^{t}z_{\tau}\geq\sum\nolimits_{\tau=1}^{t}b_{\tau}+\sum\nolimits_{\tau:\Delta_{\tau}\leq t}f_{\tau},\quad\quad\quad\quad x_{t},z_{t},s_{t}\geq 0,\quad\forall t\in[T].$
In the objective (1), $\delta\geq 0$ is a known coefficient for the delivery switching penalty, $\mathcal{S}(\cdot)$ is a function that implements the desired smoothness penalty on purchasing decisions, and $\mathcal{D}(\cdot)$ is a function that captures the delivery cost. For both of the latter terms, we consider several models inspired by real-world applications – see § 2.3 and § 2.4 for a discussion of each, respectively. We assume decisions outside of the time horizon are zero: $x_{0}=z_{0}=0$ and $x_{T+1}=z_{T+1}=0$.
As outlined above, the problem is subject to several constraints for all $t\in[T]$. In order, they are as follows: First, the storage state is governed by the inventory dynamics $s_{t}=s_{t-1}+x_{t}-z_{t}$, and the state must remain within physical limits, i.e., $0\leq s_{t}\leq S$. OSDM models a “discharged” storage state as $s_{t}=0$, but this can capture the more realistic case where $s_{t}=0$ corresponds to a minimum charge level for the storage. Next, the amount of energy purchased, i.e., $x_{t}$, must be sufficient to cover the desired delivery decision, potentially drawing from storage, but cannot overfill the storage. These constraints are captured by $z_{t}-s_{t-1}\leq x_{t}\leq z_{t}+(S-s_{t-1})$. The delivery decision at time $t$ ($z_{t}$) must cover the base demand and is allowed to satisfy any flexible demand that has arrived but is not past due – this is captured by $b_{t}\leq z_{t}\leq b_{t}+\sum_{\tau=1}^{t}f_{\tau}-\sum_{\tau:\Delta_{\tau}<t}f_{\tau}$. Finally, the cumulative delivery decisions up to $t$ must satisfy all base demand and all flexible demand whose deadlines have passed: this is captured by $\sum_{\tau=1}^{t}z_{\tau}\geq\sum_{\tau=1}^{t}b_{\tau}+\sum_{\tau:\Delta_{\tau}\leq t}f_{\tau}$. We summarize the key pieces of OSDM in Fig. 1.
While (1) captures the offline version of the problem, our goal is to design an online algorithm that chooses $\{x_{t},z_{t}\}$ at each time step $t$ without knowledge of the future online inputs of $\{p_{\tau},b_{\tau},f_{\tau},\Delta_{\tau}\}_{\tau>t}$. To do so, we follow the literature on online algorithms and competitive analysis Manasse:88; Borodin:92 and evaluate performance via the competitive ratio Manasse:88; Borodin:92, defined as follows:
@@
Note that a smaller competitive ratio implies that the online algorithm is guaranteed to be closer to the offline optimal solution. The competitive ratio is a worst-case performance metric for online algorithms; as such, bounds on this metric ensure robustness against nonstationarity in the underlying environment, which is crucial for the applications motivating OSDM. For instance, electricity prices are known to be volatile and non-stationary Qu:24, and demand profiles may be non-stationary in certain applications. However, algorithms optimized for competitive guarantees are often overly conservative in practice – to address this limitation in OSDM, we build on the learning-augmented algorithms literature Lykouris:18; Purohit:18; Mitzenmacher:22:ALPS to propose data-driven algorithms that can leverage historical data to significantly improve average-case performance while still maintaining worst-case competitive guarantees against non-stationary inputs. We discuss this further in § 6.
Figure 1: Diagram of the OSDM problem at a single (discrete) time step $t$. The demand manager specifies a purchasing decision $x_{t}$, a delivery decision $z_{t}$, and (implicitly) a storage state $s_{t}$. As a function of these decisions, they pay ➊ a purchasing cost $p_{t}x_{t}$, ➌ a smoothness penalty $\mathcal{S}(x_{t},x_{t-1})$ that discourages fluctuations in the purchasing rate, ➋ a delivery cost of serving demand $\mathcal{D}(z_{t},s_{t-1},p_{t})$, and ➍ a switching penalty $\delta|z_{t}-z_{t-1}|$ that discourages fluctuations in the delivery rate.
2.2 Motivating Applications

### v2

@@ 1. Introduction
The proliferation of renewable energy sources and grid-scale energy storage, the growth of power-hungry AI data centers, and the implementation of dynamic pricing schemes like time-of-use tariffs are fundamentally reshaping the landscape of how energy and compute infrastructures interact. These changes are prompting a paradigm shift in how systems interact with power grids, shifting from a model of passive consumption to one of active, intelligent demand management. The core challenge in demand management from an energy perspective is to better align the consumption of energy (e.g., from the grid) with external signals such as time-of-use prices, while meeting certain performance (e.g,. service-level agreement (SLA)) objectives on the demand side. Concurrently, there is a need to consider the “smoothness” of these decisions: recent reports suggest that data centers’ sudden large fluctuations in energy use are impacting power quality in U.S. grids (Nicoletti et al., 2024), and sudden changes in load are known to cause cascading failures and even blackouts on the grid, such as the 2025 Iberian Peninsula blackout (Pons and Khalip, 2025). This prompts an additional objective of smooth decisions (e.g., gradual ramping up and down).
@@
Many techniques have been developed for this class of problems, with solution designs that are usually tailored to specific applications such as energy storage management (Mamun et al., 2016) and HVAC control (Kou et al., 2021), among others. These techniques often rely on machine learning somewhere in the pipeline, e.g., to provide predictions for a model predictive controller (Serale et al., 2018; Morstyn et al., 2017), inform a day-ahead provisioning problem (Yang et al., 2025), or to directly make decisions (e.g., using a reinforcement learning agent) (Azuatalam et al., 2020). While these problem-specific solutions show promise, they are seldom deployed in practice. One reason is that environments such as electricity markets are inherently nonstationary, making it difficult to develop heuristics that deal well with distribution shift (Levin et al., 2022) or come with rigorous performance guarantees. Another reason is that these approaches are typically developed with a single end application in mind, making generalization to other settings difficult.
@@
On the other side of the spectrum, there is a rich body of work in online algorithms (Yang et al., 2020; Lechowicz et al., 2024; El-Yaniv et al., 2001; Yang et al., 2024; Chen et al., 2022) on related problems that are sufficiently abstract to model many different applications, and have been studied under frameworks such as competitive analysis, which provides rigorous performance guarantees. Despite this, however, there remains a gap between the existing models studied in the online algorithms literature and the complexity of real-world demand management problems—as of yet, no models exist that fully capture the moving pieces necessary to study the type of problem that arises in, e.g., a grid-integrated data center with local energy storage.
@@
Competitive (i.e., worst-case) guarantees are often overly pessimistic in practice, since real-world problem instances are not adversarial. In recent years, the area of learning-augmented algorithms (Lykouris and Vassilvtiskii, 2018; Kumar et al., 2018; Mitzenmacher and Vassilvitskii, 2022) has expanded the online algorithms literature with techniques that can take additional information as input (e.g., predictions about the instance) to improve average-case performance while retaining worst-case guarantees for robustness against nonstationarities. While this paradigm has successfully been applied to problems that share similarities with OSDM (Lechowicz et al., 2024; Sun et al., 2021a), traditional learning-augmented designs make strong assumptions about the availability and learnability of suitable predictions, which face challenges in OSDM due to the multiple kinds and nonstationary sources of uncertainty (e.g., demand arrivals, deadlines, and market prices). Thus, we also ask:
Can we design data-driven algorithms for OSDM that directly learn from data (e.g., historical instances) to significantly improve performance while maintaining worst-case guarantees?
1.1. Contributions
@@
In § 4, we provide the first competitive algorithm for OSDM, which we call PAAD (Partitioned Accounting & Aggregated Decisions, see Algorithm 2). PAAD introduces a driver abstraction that splits OSDM into subproblems that can be solved independently and aggregated to form a globally competitive solution (Theorem 4.3). PAAD reflects, to the best of our knowledge, the most complex extension of the online pseudo-cost minimization algorithm design paradigm (Sun et al., 2021b); we employ multiple kinds and instances of threshold functions to decompose and accommodate the significant heterogeneity of the OSDM setting. PAAD achieves a worst-case competitive ratio that is sublinear in $\nicefrac{{p_{\max}}}{{p_{\min}}}$ (the ratio of the procurement price bounds) and linear in other parameters of the problem setting such as the switching and delivery cost magnitudes—in § 5, we give lower bounds that imply PAAD achieves the best-possible competitive ratio for OSDM under our assumptions.
@@
To improve average-case performance and reduce the pessimism of our approach, in § 6 we present a learning framework for OSDM called PALD (Partitioned Accounting & Learned Decisions). Compared to existing work on learning-augmented algorithms, the OSDM setting brings unique challenges that make previously-considered techniques for integrating predictions impractical or ineffective. To address this challenge, PALD proposes a novel end-to-end differentiable learning framework that directly predicts the best algorithm parameters (e.g., threshold functions) that minimize PALD’s cost on historical instances, while ensuring that the learned algorithm satisfies a robustness certificate (see Theorem 6.3) that guarantees a bounded worst-case competitive ratio under adversarial instances, serving as a safeguard against nonstationarity. This framework will likely be of independent interest as a more flexible and expressive methodology to learn high-performing and robust algorithms across problem domains.
@@
Finally, in § 7, we present a case study of OSDM in a simulation of a grid-integrated data center with co-located battery storage, using real electricity price data and production Alibaba traces (Alibaba, 2018). We show that PAAD solves the OSDM problem effectively in practice, and that PALD’s end-to-end learning capability significantly improves performance compared to PAAD using data-driven insights.
1.2. Related Work
@@
Our study of OSDM is informed by and builds upon several lines of related work in online algorithms, learning-augmented algorithms, and relevant application areas. We briefly summarize these below.
@@
Applications of Demand Management and Related Ideas. OSDM-like problems arise across multiple domains—below, we review some heuristic and control-theoretic approaches that have been proposed to solve these. Given forecasts (e.g., prices), a common approach is to solve an offline “day-ahead” optimization problem, as applied in battery arbitrage (Mohsenian-Rad, 2016; Zhang et al., 2024; Faraji et al., 2020), microgrids (Silva et al., 2020; Tandukar et al., 2018), HVAC (Yang et al., 2025; Zeng et al., 2023), and sustainable computing (Radovanovic et al., 2022; Li et al., 2017). In online settings, a frequently-applied technique is model predictive control (MPC), which uses forecasts to solve a finite-horizon optimization problem at each time step. MPC has been applied to battery management (Blonsky et al., 2020; Chehade et al., 2024; Sage et al., 2024; Morstyn et al., 2017), building systems (Afram and Janabi-Sharifi, 2014; Serale et al., 2018; Blum et al., 2022; Mariano-Hernández et al., 2021), microgrids (Moreno-Castro et al., 2023; López-Rodríguez et al., 2021), and data centers (Wang et al., 2023; Lazic et al., 2018; Mirhoseininejad et al., 2021). Select works consider approaches that do not require predictions such as Lyapunov-based control (Urgaonkar et al., 2011; Guo and Fang, 2013; Yu et al., 2015; Stai et al., 2021; Sun et al., 2023), ADMM for distributed storage (Rigo-Mariani and Debucsshere, 2022), feedback-based demand response (Mamun et al., 2016), and stochastic optimization (Nasiriani et al., 2017; Bedi et al., 2018). These approaches are generally developed with a single application in mind, limiting their generalizability, and do not typically come with theoretical guarantees on performance, particularly in cases where predictions or stochastic models are used.
@@
Related Online Algorithms Problems. In the literature there are a few online problems that share structure with OSDM. Of these, perhaps the closest related problem is online linear inventory management (OLIM(Yang et al., 2020), which models the problem faced by an operator (e.g., a data center) with an inventory (e.g., battery) that must meet a demand at each time step, while minimizing their procurement cost subject to time-varying prices. Compared to OLIM, OSDM additionally models a delivery cost to serve demand (i.e., separate from procurement), smoothing penalties, and a mixture of base (inelastic) and flexible demand. A recent model called online conversion with switching costs (OCS(Lechowicz et al., 2024) has modeled the problem faced by an operator procuring energy (e.g., an EV charger) that must meet a single flexible demand before a deadline (e.g., charging an EV by the end of the day), paying a switching cost when decisions change. OSDM generalizes this by including inflexible (i.e., base) demand constraints, delivery costs, and multiple flexible units with heterogeneous deadlines. Other formulations address energy storage without time-varying procurement prices (Kim et al., 2016; Mo et al., 2021) or microgrid scheduling with variable prices but no storage (Lu et al., 2013; Hajiesmaili et al., 2016; Zhang et al., 2018).
@@
Each component dynamic of OSDM has been studied in isolation: procurement and flexible demand via online knapsack and search (Zhou et al., 2008; El-Yaniv et al., 2001; Fung, 2021), per-step demand via online matching (Mehta et al., 2007; Huang et al., 2024), and smoothness penalities via metrical task systems (Borodin et al., 1992; Friedman and Linial, 1993) However, the combination of these in OSDM introduces new technical challenges in the design of robust and efficient algorithms that preclude the direct application of existing results (see, e.g., § 3). Online resource allocation problems (Jaillet and Lu, 2011; Devanur et al., 2019; Wang et al., 2017; Jiao et al., 2017; Klimm et al., 2019; Zhang et al., 2017; Balseiro et al., 2023) are also related, though they usually assume a fixed (non-replenishable) resource. Some variants consider inventories (Larsen and Wøhlk, 2010; Cheung et al., 2021; Chen et al., 2022; He et al., 2022; Hihat et al., 2023; Yang et al., 2024), but typically treat either demand or inventory evolution as exogenous, whereas in OSDM both are decisions.
2. Problem Formulation
@@
2.1. Online Smoothed Demand Management (OSDM)
Problem Statement. Consider an operator who makes decisions over a discrete time horizon $t\in[T]$. At each time step $t$, several inputs are revealed online. The first of these is a market price for energy, denoted by $p_{t}\geq 0$. Furthermore, two types of demand arrive: a base (inflexible) demand $b_{t}\geq 0$ that must be satisfied immediately, and a flexible demand $f_{t}\geq 0$ that can be deferred until a specified deadline $\Delta_{t}>t$. Further, the operator can control the state of a local energy storage unit (e.g., a battery) with a maximum capacity of $S\geq 0$. We denote the state of charge at time step $t$ as $s_{t}$, where the initial storage state is empty (i.e., $s_{0}=0$). [^1] The assumption that the initial storage state is $s_{0}=0$ is without loss of generality—our results hold when $s_{0}>0$.
@@ 2.1. Online Smoothed Demand Management (OSDM)
(1) $[\texttt{OSDM}]\min_{\{x_{t},z_{t}\}_{t\in[T]}}$ $\sum\nolimits_{t=1}^{T}\Big[\underbrace{p_{t}x_{t}}_{\text{purchase cost}}+\underbrace{\mathcal{D}(z_{t},s_{t-1},p_{t})}_{\text{delivery cost}}\Big]+\sum\nolimits_{t=1}^{T+1}\Big[\underbrace{\mathcal{S}(x_{t},x_{t-1})+\delta|z_{t}-z_{t-1}|}_{\text{smoothing penalty}}\Big],$
s.t. ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{t}=s_{t-1}+x_{t}-z_{t},\text{ with }0\leq s_{t}\leq S,}$
@@
(2) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum\nolimits_{t=\tau_{1}}^{\tau_{2}}z_{t}-b_{t}\geq\sum\nolimits_{t:t\geq\tau_{1}\text{ and }\Delta_{t}\leq\tau_{2}}f_{t}\quad\quad\forall(\tau_{1},\tau_{2}):\tau_{1}\leq\tau_{2}\text{ and }\tau_{1},\tau_{2}\in[T].}$
In the objective (1), $\delta\geq 0$ is a known coefficient for the delivery switching penalty, $\mathcal{S}(\cdot)$ is a function that implements the desired smoothness penalty on purchasing decisions, and $\mathcal{D}(\cdot)$ is a function that captures the delivery cost. For both of the latter terms, we consider several models inspired by real-world applications—see § 2.3 and § 2.4 for a discussion of each, respectively. We assume decisions outside of the time horizon are zero: $x_{0}=z_{0}=0$ and $x_{T+1}=z_{T+1}=0$.
As outlined above, the problem is subject to several constraints for all $t\in[T]$. In order, they are as follows: First, the storage state is governed by the inventory dynamics $s_{t}=s_{t-1}+x_{t}-z_{t}$, and the state must remain within physical limits, i.e., $0\leq s_{t}\leq S$. OSDM models a “discharged” storage state as $s_{t}=0$, but this can capture the more realistic case where $s_{t}=0$ corresponds to a minimum charge level for the storage. Next, the amount of energy purchased, i.e., $x_{t}$, must be sufficient to cover the desired delivery decision, potentially drawing from storage, but cannot overfill the storage. These constraints are captured by $z_{t}-s_{t-1}\leq x_{t}\leq z_{t}+(S-s_{t-1})$. The delivery decision at time $t$ ($z_{t}$) must cover the base demand and is allowed to satisfy any flexible demand that has arrived but is not past due—this is captured by $b_{t}\leq z_{t}\leq b_{t}+\sum_{\tau=1}^{t}f_{\tau}-\sum_{\tau:\Delta_{\tau}<t}f_{\tau}$. Finally, the cumulative delivery for any subinterval of $[T]$ must satisfy base demand and any flexible demands whose lifetime lies within this interval: this is captured by $\sum\nolimits_{t=\tau_{1}}^{\tau_{2}}z_{t}-b_{t}\geq\sum\nolimits_{t:t\geq\tau_{1}\text{ and }\Delta_{t}\leq\tau_{2}}f_{t}\;\forall(\tau_{1},\tau_{2}):\tau_{1}\leq\tau_{2}\text{ and }\tau_{1},\tau_{2}\in[T]$. We summarize the key pieces of OSDM in Fig. 1. We remark that if flexible demands are not considered, the two decision variables in the OSDM formulation can be simplified to a single decision variable (i.e., $\{x_{t}-z_{t}\}_{t\in[T]}$)—this simplified version of OSDM is of independent interest, and we refer to it as “OSDM with only base demand” throughout the paper (see e.g., Corollary 5.2 for tighter results in this regime).
While (1) captures the offline version of the problem, our goal is to design an online algorithm that chooses $\{x_{t},z_{t}\}$ at each time step $t$ without knowledge of the future online inputs of $\{p_{\tau},b_{\tau},f_{\tau},\Delta_{\tau}\}_{\tau>t}$. To do so, we follow the literature on online algorithms and competitive analysis (Manasse et al., 1988; Borodin et al., 1992) and evaluate performance via the competitive ratio (Manasse et al., 1988; Borodin et al., 1992), defined as follows:
@@
Note that a smaller competitive ratio implies that the online algorithm is guaranteed to be closer to the offline optimal solution. The competitive ratio is a worst-case performance metric for online algorithms; as such, bounds on this metric ensure robustness against nonstationarity in the underlying environment, which is crucial for the applications motivating OSDM. For instance, electricity prices are known to be volatile and non-stationary (Qu et al., 2024), and demand profiles may be non-stationary in certain applications. However, algorithms optimized for competitive guarantees are often overly conservative in practice—to address this limitation in OSDM, we build on the learning-augmented algorithms literature (Lykouris and Vassilvtiskii, 2018; Kumar et al., 2018; Mitzenmacher and Vassilvitskii, 2022) to propose data-driven algorithms that can leverage historical data to significantly improve average-case performance while still maintaining worst-case competitive guarantees against non-stationary inputs. We discuss this further in § 6.
Figure 1. Diagram of the OSDM problem at a single (discrete) time step $t$. The demand manager specifies a purchasing decision $x_{t}$, a delivery decision $z_{t}$, and (implicitly) a storage state $s_{t}$. As a function of these decisions, they pay ➊ a purchasing cost $p_{t}x_{t}$, ➌ a smoothness penalty $\mathcal{S}(x_{t},x_{t-1})$ that discourages fluctuations in the purchasing rate, ➋ a delivery cost of serving demand $\mathcal{D}(z_{t},s_{t-1},p_{t})$, and ➍ a switching penalty $\delta|z_{t}-z_{t-1}|$ that discourages fluctuations in the delivery rate.
2.2. Motivating Applications